OpenStudy (anonymous):

simplify -14x^3/x^3-5x^4

OpenStudy (anonymous):

So, do you know how to do any of this?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Hm, ok then.

OpenStudy (anonymous):

The first thing you have to do is combine like terms, now, every thing in here are like terms, and we can do things to X as long as we are multiplying or dividing.

OpenStudy (anonymous):

-14x^3/x^3-5x^4 So, -14 to the 3rd power yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we have to mulitply it to the 3rd power, making 14 * 14 is 192, 14

OpenStudy (anonymous):

14* 14= 192, * 14= 778 1920

OpenStudy (anonymous):

x2698

OpenStudy (anonymous):

thats -14 to the 3rd power in simples form(-x2698 , now we divide by the next one

OpenStudy (anonymous):

sorry its taking me so long, I try not to use a calculator, any how -14x^3/x^3-5x^4 do you know what x to the 3rd power is?

OpenStudy (anonymous):

X is just 1

OpenStudy (anonymous):

so you did 14*14*14 and go 2698

OpenStudy (anonymous):

mm hmmm

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

isn't it 2744

OpenStudy (anonymous):

yes, thank you, like I said I don't use calculators.

OpenStudy (anonymous):

ok not confused now lol

OpenStudy (anonymous):

ok then, lets keep going, x to the 3rd power is just 1, so 1 divided by 2744 is -x2744.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-14x^3/x^3-5x^4 now then ... it is Negative 5x or minus?

OpenStudy (anonymous):

-5x to the 4th power = 5,25, 125, 625

OpenStudy (anonymous):

-x625-2744=-x2118

OpenStudy (anonymous):

NOW

OpenStudy (anonymous):

you have to understand, that because this problems variables were all completely togherter and had different methods such as subtracting dividing, this must have actually already been in simplest form

OpenStudy (anonymous):

Otherwise, in problems like -14x^3(3) /x^3-5x^4 you would of had to mulitply everything and add.

OpenStudy (anonymous):

already then, im glad I was of help.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are very welcome.